# EgoBlind: Towards Egocentric Visual Assistance for the Blind People

### v1

@@ 4.2.1 Overall Observations
Table 5 presents the performances of different models across different question types. We outline the following main observations below:
@@
  • The closed-source models perform similarly to the top-performing open-source ones such as InternVL2 [5] and LLaVA-OV [19].
  • The models that are superior at general-purpose ego-VQA and image blind-QA are not the best-performing.
  • No single model wins across all question types.
  • Answering “Navigation” questions is the most challenging task for almost all models, followed by “Tool Use”.
  • Better LLMs always bring but are not clear-cut for better performance (Qwen2 vs. LLaMA3 vs. Vicuna).
  • Large variant of the same models (e.g., InternVL2-26B vs. 8B) may not necessarily be the stronger one.
@@ 4.2.2 Specific Challenges
Figure 5: MLLMs face challenges in understanding user intention. The question is posed at the last frame (same as below).
User Intention. We found that, even when the prompt emphasizes that the model should serve blind people, the models still fail to fully understand the context and intent of the question in real scenarios, leading to inaccuracies in its answers. For a typical example in Fig. 5, where a blind person asks, “Is there a road ahead?” although a road is visible in the image, the subject’s path is actually blocked by a number of obstacles, such as yellow-black roadblocks, lamp posts, and large trees. A reasonable response should be “No,” with further recommendations for an alternative path, as suggested by all the ground-truth answers. However, the model generally tends to answer “Yes”, without considering the impact of these obstacles for navigation. For improvements, it is crucial to spatio-temporally engage into the users’ real-time activities.
Figure 6: MLLMs face challenges in reasoning spatial orientation.
Real-time Spatial Orientation. MLLMs exhibit significant shortcomings in understanding spatial orientations relative to the subjects (blind people) when they are moving. A typical example is shown in Fig. 6. Specifically, when the subject moves, the models fail to effectively update their real-time spatial orientations relative to the subject (e.g., “right side”), and favor static and prominent objects (e.g., “oil”) while overlooking the actual needs from the subjecs’ perspectives (e.g., “tomato”). A similar case is also found in Fig. 8, where all models (except for Gemini 2.0) fail to assist the user in localizing the escalator. This posits to a limitation towards egocentric embodied AI assistance. For improvements, the models should be sensitive to the subtle visual difference in continuous video frames.
Temporal Context. MLLMs often fail to reason the temporal contexts to precisely understand and answer the question. For example in Fig. 7, the blind user asks a simple question “where is my companion?”. Only Gemini gets close responses, possibly because of its raw video inputs (continuous frames and audio). Most models fail to correctly answer the question. It seems that the models do not know who is the companion of the blind user, and hallucinate their answers that mismatch the visual fact (e.g., “on the boat”, “on the left”). For human, it is easy to know that the “man in grey” is the user’s companion even by a simple glance over the video content. Similarly, Fig. 8 demonstrates that most models fail to recall the temporal moment (the 1st frame) that shows the location of the escalator. This discrepancy reveals a model limitation in reasoning the temporal contexts. Long-memory techniques could help address this limitation.
Figure 7: MLLMs fail to reason the temporal contexts.
Figure 8: MLLMs fail to reason the spatio-temporal contexts.
Identifying Obstacle. MLLMs’ performance in “safety warning” questions remains inadequate. They cannot serve as reliable safety tools to provide users with accurate answers. For example, in a typical scenario where the subject is walking along the road and asks “Are there any obstacles in my path?” the answers generated by the models often show significant deviations from the real situation (Ground Truth, GT), exhibiting a mismatch with the question asked. Even though there may be potential obstacles in the scene (such as pedestrians, vehicles, or other barriers), the models fail to accurately identify and alert the user, instead providing irrelevant or vague responses. This discrepancy not only affects the accuracy of the answers but also reflects a mismatch between the style of the answers and the user’s actual needs. Despite having strong scene detection abilities, the models still fall short in understanding the user’s intent within context, identifying potential risks, and providing precise warnings. For improvement, we hope our training data can serve as a valuable resource.
Figure 9: MLLMs face challenges in identifying obstacle.
Figure 10: MLLMs tend to be sycophant.
Sycophancy. MLLMs exhibit significant shortcomings when the blind users ask non-existent object due to blindness. For example, in a typical scenario where a blind person, sitting in a car with no other passengers, asks “Who is sitting next to me?”, only human could correctly understand the intent and provide a negative answer (e.g., “No one is sitting next to you”). In contrast, the models often deviate from the actual visual situation or the question. For instance, some models hallucinate that there are other passengers in the car, while others identify the driver as sitting next to the user but fail to give a negative answer. This sycophantic phenomenon highlights the limitations of current MLLMs when dealing with unconventional questions from blind users. Due to their visual impairment, blind individuals often ask questions that may seem “strange” to sighted individuals but are entirely reasonable from their perspective. For alleviation, synthesized data for tailored training could be helpful.
@@ 5 Conclusion
In this paper, we present the construction of a VideoQA dataset EgoBlind to reflect on the challenges towards egocentric visual assistance for the blind people. EgoBlind is the first of its kind in that the videos are token by the visually-impaired individuals from a first-person perspective, with the questions posed and verified by blind users to ensure close alignment with their true assistant needs. We further comprehensively benchmark the challenge with 15 prominent multimodal LLMs and unveil their significant limitations in various aspects. We conduct thorough analysis and share many key insights for future advancements in this direction. By formulating and bringing the challenge to the vision-language community, our primary goal is to push the MLLM research towards live egocentric visual assistance for the overwhelming number of visually impaired people around the world.
Importantly, we have ensured that the dataset collection and user studies adhere to IRB standards. Informed consent was obtained from all the content creators, who agreed to their content being used for non-commercial research purposes. Access to the dataset will be granted via authorized signatories, and video sources will be cited for distribution.
@@ Appendix A EgoBlind Dataset Creation
A.1 AI Generation
The process of calling the GPT-4o API to generate QA pairs is mainly divided into three parts. Video Decoding and Frame Extraction. We first perform video abstraction on the edited video clip by saving one frame every 10 frames. The process of generating QAs from video involves several steps: The video is decoded into individual frames. For efficiency, every 10th frame is extracted (saved), which reduces the amount of data while retaining key visual information. Uniform Sampling. After initial frame extraction, further uniform sampling is applied. The sampling interval varies based on the video length to balance resource usage and data redundancy. For instance: A 3-minute video might be sampled every 18 frames. A 30-second video might be sampled every 6 frames. This adaptive approach ensures that shorter videos have more frequent samples, capturing essential details, while longer videos maintain a broader overview. GPT-4o Model Input The uniformly sampled frames are input into the GPT-4o model. Each frame is accompanied by its timestamp and a refined prompt (as shown in LABEL:tab:gen_qa), instructing the model to generate relevant, real-time question-answer pairs based on the frame content. Batch Processing for Diversity. To enhance data diversity, the generation process is divided into batches by question categories. This means that each batch contains only one type of question, which can help in organizing and enriching the dataset, ensuring a wide coverage of different topics and improving overall quality. This structured approach ensures that the generated QA pairs are both relevant and diverse, effectively utilizing video content to create a comprehensive and organized dataset. The specific prompts for each category are attached in LABEL:tab:gen_qa.
@@ A.2 Blind User Study
We categorize manually corrected QA pairs (translated into Chinese) based on video scenes and invite 111 blind individuals to evaluate the blind-relevance of the questions across different scenarios. The evaluation scale ranges from 0 to 5, with higher scores indicating preferable questions and lower scores reflecting unwanted questions (as illustrated in Fig. 13, a real example is shown in Fig. 20). Additionally, blind individuals are encouraged to provide further questions and insights to better meet the needs of individuals with visual impairments.
Figure 11: Distribution of blind participants.
The cohort study (Fig. 11) exhibits substantial demographic heterogeneity, encompassing participants ranging from teenagers to older adults (Fig. 12). Male participants account for a dominant proportion of 73.3%, with young adults aged 18-30 constituting the largest age cohort. Notably, for most respondents, the duration of visual impairment corresponded with their biological age, suggesting congenital origin - a finding consistent with the observation that over 80% presented Grade 1 Blindness (total visual acuity $\leq$0.02 or visual field <5absent5<5< 5°). The remaining participants acquire visual impairment postnatally, though all reported living with the condition for multiple years.
@@
This demographic and geographic diversity ensures comprehensive coverage of life scenes across varying urban contexts, from high-density metropolitan transit networks to community-level infrastructure systems.
Figure 12: Distribution of user ages and their blind ages.
Figure 13: The scoring example
@@
We found that when blind people visit supermarkets, stores, and shopping malls (as shown in Fig. 14); their primary concerns are navigation and positioning issues related to location. Specifically, they want to know where they are and how to get to their desired destination(“How to get to XXX from here?”) . Their second main concern is obtaining product information, such as price (“Where can I find the price of this product?”), style, shelf life, etc. Additionally, some blind individuals are also interested in understanding the range of products sold at specific stores and the layout of the mall, including questions like “What does this store sell?” and “Which products are displayed on the counter?” At the same time, we found that blind individuals tend to be less concerned about the presence and location of people around them. Questions related to the number of people nearby (“How many people are in front of me?”) and their specific whereabouts (“Is the cashier directly in front of me?”) received a significantly higher proportion of low scores compared to high scores. Additionally, they showed minimal interest in information about road obstacles. The proportion of low scores for questions such as ”What are the obstacles on the road?” and ”How many floors are there?” was considerably higher than the proportion of high scores.
@@
Figure 14: MOS score distribution of blind people in shopping scenarios. Each row of the ordinate represents a question, and the stacked parts of each color represent the distribution proportions of different scores.
Figure 15: MOS score distribution of blind people using transportation. Each row of the ordinate represents a question, and the stacked parts of each color represent the distribution proportions of different scores.
@@
Figure 16: MOS score distribution of blind people navigating outdoor environments. Each row of the ordinate represents a question, and the stacked parts of each color represent the distribution proportions of different scores.
Figure 17: MOS score distribution of blind people in shopping scenarios. Each row of the ordinate represents a question, and the stacked parts of each color represent the distribution proportions of different scores.
@@
Figure 18: MOS score distribution of blind people in cooking and doing craft scenarios. Each row of the ordinate represents a question, and the stacked parts of each color represent the distribution proportions of different scores.
@@
When blind people use transportation such as cars, buses, and other vehicles (as shown in Fig. 15), over 60% of the respondents assigned full scores to questions like ”Is this my taxi?” and ”Which bus is coming?” while waiting for a taxi. This indicates that their primary concern is the current location of the vehicle. Meanwhile, respondents uniformly gave low scores to the question ”Is the car being moved?” suggesting that blind individuals are capable of perceiving vehicle movement. Furthermore, approximately 80% of respondents gave low scores to questions such as ”What does the bus stop advertisement say?” and ”What brand is the car?” indicating that blind people often show little interest in irrelevant information surrounding their waiting position.
@@
In outdoor navigation scenarios, such as finding destinations, crossing roads, or walking, blind individuals prioritize safety above all else (as shown in Fig. 16). For instance, the question “Is it safe to cross the road?” receives the highest score of 5 from all respondents. Similarly, over 80% of respondents rate questions related to safety during movement, such as “How can I safely get through this entrance?” and “Is it safe to go straight?”, with scores of 4 or higher. Blind individuals also place significant emphasis on directional and positional information. For questions about accessing public facilities, such as “How to reach the nearest sidewalk, zebra crossing, or blind road?”, over 70% of respondents gave scores of 4 or higher. Direction-related questions like “Which direction should I take to reach the intersection?” and “Is the supermarket on my left?” received scores above 3 from more than 60% of respondents. However, respondents show less concern for questions about the characteristics of the surrounding environment, such as “Please describe the environment around me” and “Is traffic busy in this area?”. Similarly, questions about weather conditions (“Is it sunny now?”) or information about nearby pedestrians (“Who is the person standing in front?”, “Are there many people around us gathered?”, “Is there a pedestrian coming towards me?”) received scores below 3 from more than 50% of respondents, indicating a general lack of interest. Interestingly, questions about obstacles did not receive as high a score as might be intuitively expected. For example, fewer than 50% of respondents positively rated questions like “Can you describe the obstacles ahead?” and “To avoid obstacles, which direction should I go?”, while a large proportion gave neutral scores (3). This may be due to differences in the proficiency of using white canes among blind individuals, which influences their reliance on such information. Similarly, questions such as “Am I walking on a blind road” also exhibited a polarized response, with respondents either assigning very low scores (0 or 1) or very high scores (4 or 5). This polarization suggests that reliance on certain types of information may vary significantly depending on individual preferences and skills.
@@
We also observed characteristics of blind individuals’ activities in familiar environments that contradict common assumptions. As shown in Fig. 15, when blind individuals are at home, we found that the proportion of questions receiving low scores (0, 1, or 2) exceeded 60% across almost all queries. Respondents explained that although they cannot see, their brains form a mental “ma” of the environment based on prior familiarity. This mental representation allows them to navigate and plan their movements seamlessly within the environment. Unless there are significant changes in the environment, they typically do not encounter any issues.
@@
To validate the rationality of question design in hands-on activity scenarios for visually impaired individuals, we conducted a study where participants rated questions in scenarios such as cooking and crafting (As shown in Fig. 18). Taking these two scenarios as examples, we found that, despite the differences in context, respondents consistently gave high scores (4 and 5) to core questions that directly influence task completion. For instance, “How much time is left?” (Cooking) and “Please describe the items on the shelves” (Doing crafts). At the same time, due to the differences in activities, visually impaired individuals tended to give a higher proportion of low scores (0 or 1) to cooking-related questions compared to crafting. This can be attributed to the fact that cooking typically occurs in familiar environments where blind individuals rely on their spatial memory and experience. In contrast, in the crafting scenario, respondents showed a neutral or low level of interest (0–3) in questions about basic item information such as appearance or shape, exemplified by questions like “Is it a cup?” and “What does this cup look like?”.
Appendix B Experimental

### v2

@@ 4.2.1 Overall Observations
Table 4 presents the performances of different models. We summarize the following observations:
@@
  • The models that are superior at general-purpose egocentric VQA (e.g., LLaVA-Video) and image blind-VQA (e.g., VILA1.5) are not the best-performing, depicting unique challenges of EgoBlind.
  • No single model wins across all question types. Answering “Navigation” questions is the most challenging task for almost all models, indicating a significant limitation of MLLMs in this field.
  • While most other models struggle in answering questions about tool use, Gemini 1.5 even surpasses human performance, demonstrating its rich knowledge outweighing human individuals.
  • Stronger LLMs and larger visual resolution often bring better performance, while more frames do not always help (e.g., 8 frames are enough for InternVL to surpass other open-source models.).
@@ 4.2.2 Specific Challenges
We further reveal some specific challenges pertaining to egocentric visual assistance for the blind by analyzing the common failure cases.
Figure 4: Common failure cases of tested MLLMs. The models fail to (a) reason user intention, (b) understand real-time spatial orientation, provide (c) assistive and (d) reliable answers.
User Intention. Reasoning user intentions behind the questions is key for effective assistance. Yet, all models fall short in such capability. For example, in Figure 4(a) and (c), all models fail to generate helpful answers for the blind, though the answers are objectively correct to the visual contents. It is worth mentioning that we have explicitly prompted the models to answer questions of blind individuals to provide visual assistance.
Spatio-temporal Orientation. MLLMs exhibit significant shortcomings in weaving together the temporal frames to reason the spatial orientations relative to the users’ real-time location. A typical example is shown in Figure 4(b), where indoor navigation is requested by the blind user. The escalator was framed prior to the question moment, and the models have to retrieve the related moment and reason the users’ orientation change after that moment, which shows extreme challenge to all MLLMs.
Reliable Answers. Providing reliable answers is of crucial importance in egocentric visual assistance. However, all MLLMs tend to be sycophantic when the users ask something that deviates the visual facts due to blindness. Take Figure 4(d) as an example, the blind user is already on the bridge but ask for navigating it. All models fail to remind the user that he is on the bridge, but either answer “I do not know”, or give wrong or even malicious suggestions, such as “take the boat”, “walk straight”.
We also analyze other challenges related to streaming VQA, scene text recognition, and referential words in spoken languages in the Appendix B.2.
Table 5: Single frame inputs.
Models Acc. (bef) Acc. (aft)
InternVL2.5-26B 56.6 55.5 $\downarrow$ 1.1
VILA1.5 47.2 45.8 $\downarrow$ 1.4
GPT-4o 60.6 60.2 $\downarrow$ 0.4
Table 6: Normal QA prompts.
Models Acc. (bef) Acc. (aft)
InternVL2.5-26B 56.6 54.7 $\downarrow$ 1.9
Gemini 1.5 Flash 57.9 52.5 $\downarrow$ 5.4
GPT-4o 60.6 56.8 $\downarrow$ 3.8
Table 7: Instruction Tuning.
Models Acc. (bef) Acc. (aft)
Qwen2.5-VL 49.0 53.5 $\uparrow$ 4.5
LLaVA-OV 55.1 57.5 $\uparrow$ 2.4
InternVL2.5-8B 55.8 62.1 $\uparrow$ 6.3
4.3 Other Investigations
Our investigations aim to answer three questions: 1) Is a single frame at the question timestamp enough to answer the question? 3) Whether it is helpful to explicitly prompt the models that the questions are asked by blind users for visual assistance? 2) Whether using our training data for instruction tuning can improve the performance?
Table 4.2.2 shows that replacing multi-frame inputs with a single frame at the question timestamp degenerates model performances, though a single frame brings competitive results (which is reasonable as people often ask about their current visual environment.). Table 4.2.2 shows that substituting the blind-specific prompt with a normal VQA prompt significantly jeopardize the model performance. This reveals that answering questions from blind people sparks new challenges compared with QA about general visual understanding. Table 4.2.2 shows that using our training data for instruction tuning (LoRA finetune) can remarkably improve the model performances, demonstrating the importance of our training data towards improvements. We present other details in Appendix B.3.
@@ 5 Conclusion
We present the construction of a VideoQA dataset EgoBlind to reflect on the challenges towards egocentric visual assistance for the blind people. EgoBlind is the first of its kind in that the videos are token by the visually-impaired individuals from a first-person perspective, with the questions posed and verified by blind users to ensure close alignment with their true assistant needs. We further comprehensively benchmark the challenge with multiple prominent multimodal LLMs and unveil their significant limitations in various aspects. We conduct thorough analysis and share many key insights for future advancements in this direction. By formulating and bringing the challenge to the vision-language community, our primary goal is to push the MLLM research towards live egocentric visual assistance for the overwhelming number of visually impaired people around the world.
Importantly, we have ensured that the dataset collection and user studies adhere to IRB standards. Informed consent was obtained from all the content creators, who agreed to their content being used for non-commercial research purposes. Access to the dataset will be granted with the condition that video sources will be cited for distribution.
@@ Appendix A EgoBlind Dataset Creation
A.1 QA Generation with GPT-4o
To enrich the QA diversity and reduce the annotation burden, we prompt GPT-4o to act as blind users to generate part of questions and invite both blind and sighted people to check and edit them. For generation, we first decode the video into frames at 3 fps. We then adopt an adaptive sampling strategy over the decoded frames to obtain the final frames to be fed to GPT-4o for QA generation. We sample at a ratio of every 18 frames for long videos (e.g., 180s) and 6 frames for short ones (e.g., 30s). To simulate real-time QA generation, each frame is accompanied by its timestamp and a refined prompt (shown in Table LABEL:tab:gen_qa). The generation process is divided into batches by question categories. This means that each batch contains only one type of question, which can help in organizing and enriching the dataset, ensuring a wide coverage of different topics and improving overall quality. The specific prompts for each category are attached in Table LABEL:tab:gen_qa.
@@ A.2 Blind User Study
We categorize manually corrected QA pairs (translated into Chinese) based on video scenes and invite 111 blind individuals to evaluate the blind-relevance of the questions across different scenarios. The evaluation scale ranges from 0 to 5, with higher scores indicating preferable questions and lower scores reflecting unwanted questions (as illustrated in Figure 5, a real example is shown in Figure 19). Additionally, blind individuals are encouraged to provide further questions and insights to better meet the needs of individuals with visual impairments.
Figure 5: The scoring example
The cohort study (Figure 6(a)) exhibits substantial demographic heterogeneity, encompassing participants ranging from teenagers to older adults (Figure 6(b)). Male participants account for a dominant proportion of 73.3%, with young adults aged 18-30 constituting the largest age cohort. Notably, for most respondents, the duration of visual impairment corresponded with their biological age, suggesting congenital origin - a finding consistent with the observation that over 80% presented Grade 1 Blindness (total visual acuity $\leq$0.02 or visual field <5absent5<5< 5°). The remaining participants acquire visual impairment postnatally, though all reported living with the condition for long time.
(a) Location distribution.
(b) Age distribution.
Figure 6: Distribution of blind participants
@@
This demographic and geographic diversity ensures comprehensive coverage of life scenes across varying urban contexts, from high-density metropolitan networks to community-level systems.
@@
We found that when blind people visit supermarkets, stores, and shopping malls (as shown in Figure 7); their primary concerns are navigation and positioning issues related to location. Specifically, they want to know where they are and how to get to their desired destination(“How to get to XXX from here?") . Their second main concern is obtaining product information, such as price (“Where can I find the price of this product?"), style, shelf life, etc. Additionally, some blind individuals are also interested in understanding the range of products sold at specific stores and the layout of the mall, including questions like “What does this store sell?" and “Which products are displayed on the counter?" At the same time, we found that blind individuals tend to be less concerned about the presence and location of people around them. Questions related to the number of people nearby (“How many people are in front of me?") and their specific whereabouts (“Is the cashier directly in front of me?") received a significantly higher proportion of low scores compared to high scores. Additionally, they showed minimal interest in information about road obstacles. The proportion of low scores for questions such as "What are the obstacles on the road?" and "How many floors are there?" was considerably higher than the proportion of high scores.
@@
Figure 7: MOS score distribution of blind people in shopping scenarios. The stacked parts of each color represent the distribution proportions of different scores (same for below).
Figure 8: MOS score distribution of blind people using transportation.
@@
Figure 9: MOS score distribution of blind people navigating outdoor environments.
Figure 10: MOS score distribution of blind people in shopping scenarios.
@@
Figure 11: MOS score distribution of blind people in cooking and doing craft scenarios.
@@
When blind people use transportation such as cars, buses, and other vehicles (as shown in Figure 8), over 60% of the respondents assigned full scores to questions like "Is this my taxi?" and "Which bus is coming?" while waiting for a taxi. This indicates that their primary concern is the current location of the vehicle. Meanwhile, respondents uniformly gave low scores to the question "Is the car being moved?" suggesting that blind individuals are capable of perceiving vehicle movement. Furthermore, approximately 80% of respondents gave low scores to questions such as "What does the bus stop advertisement say?" and "What brand is the car?" indicating that blind people often show little interest in irrelevant information surrounding their waiting position.
@@
In outdoor navigation scenarios, such as finding destinations, crossing roads, or walking, blind individuals prioritize safety above all else (as shown in Figure 9). For instance, the question “Is it safe to cross the road?" receives the highest score of 5 from all respondents. Similarly, over 80% of respondents rate questions related to safety during movement, such as “How can I safely get through this entrance?" and “Is it safe to go straight?", with scores of 4 or higher. Blind individuals also place significant emphasis on directional and positional information. For questions about accessing public facilities, such as “How to reach the nearest sidewalk, zebra crossing, or blind road?", over 70% of respondents gave scores of 4 or higher. Direction-related questions like “Which direction should I take to reach the intersection?" and “Is the supermarket on my left?" received scores above 3 from more than 60% of respondents. However, respondents show less concern for questions about the characteristics of the surrounding environment, such as “Please describe the environment around me" and “Is traffic busy in this area?". Similarly, questions about weather conditions (“Is it sunny now?") or information about nearby pedestrians (“Who is the person standing in front?", “Are there many people around us gathered?", “Is there a pedestrian coming towards me?") received scores below 3 from more than 50% of respondents, indicating a general lack of interest. Interestingly, questions about obstacles did not receive as high a score as might be intuitively expected. For example, fewer than 50% of respondents positively rated questions like “Can you describe the obstacles ahead?" and “To avoid obstacles, which direction should I go?", while a large proportion gave neutral scores (3). This may be due to differences in the proficiency of using white canes among blind individuals, which influences their reliance on such information. Similarly, questions such as “Am I walking on a blind road” also exhibited a polarized response, with respondents either assigning very low scores (0 or 1) or very high scores (4 or 5). This polarization suggests that reliance on certain types of information may vary significantly depending on individual preferences and skills.
@@
We also observed characteristics of blind individuals’ activities in familiar environments that contradict common assumptions. As shown in Figure 8, when blind individuals are at home, we found that the proportion of questions receiving low scores (0, 1, or 2) exceeded 60% across almost all queries. Respondents explained that although they cannot see, their brains form a mental “ma” of the environment based on prior familiarity. This mental representation allows them to navigate and plan their movements seamlessly within the environment. Unless there are significant changes in the environment, they typically do not encounter any issues.
@@
To validate the rationality of question design in hands-on activity scenarios for visually impaired individuals, we conducted a study where participants rated questions in scenarios such as cooking and crafting (As shown in Figure 11). Taking these two scenarios as examples, we found that, despite the differences in context, respondents consistently gave high scores (4 and 5) to core questions that directly influence task completion. For instance, “How much time is left?” (Cooking) and “Please describe the items on the shelves” (Doing crafts). At the same time, due to the differences in activities, visually impaired individuals tended to give a higher proportion of low scores (0 or 1) to cooking-related questions compared to crafting. This can be attributed to the fact that cooking typically occurs in familiar environments where blind individuals rely on their spatial memory and experience. In contrast, in the crafting scenario, respondents showed a neutral or low level of interest (0–3) in questions about basic item information such as appearance or shape, exemplified by questions like “Is it a cup?” and “What does this cup look like?”.
A.3 Analysis of Videos
We analyze more details about the videos in EgoBlind. First, unlike the famous Ego4D [32] videos which are framed in a top-down view of sighted people via head-mounted camera devices (e.g., smart glasses), the videos in EgoBlind are captured by GoPros mounted in front of the blind users’ chests or mobile phones held in their hands. Specific challenges resulting from EgoBlind videos are: 1) The objects of interest are often off-center and out-of-focus due to blindness, as shown in Figure 12(a), 2) the objects are shown in challenging viewpoint (blind users tend to walk alongside a street curb or a building for safety.), and 3) The scene could be largely occluded by their canes, mobile phones, guide dogs or other people.
Figure 12: Visualization of EgoBlind examples.
Appendix B Experiment